# SDSC:A Structure-Aware Metric for Semantic Signal Representation Learning

### v1

@@ Abstract
We propose the Signal Dice Similarity Coefficient (SDSC), a structure-aware metric function for time series self-supervised representation learning. Most Self-Supervised Learning (SSL) methods for signals commonly adopt distance-based objectives such as mean squared error (MSE), which are sensitive to amplitude, invariant to waveform polarity, and unbounded in scale. These properties hinder semantic alignment and reduce interpretability. SDSC addresses this by quantifying structural agreement between temporal signals based on the intersection of signed amplitudes, derived from the Dice Similarity Coefficient (DSC).Although SDSC is defined as a structure-aware metric, it can be used as a loss by subtracting from 1 and applying a differentiable approximation of the Heaviside function for gradient-based optimization. A hybrid loss formulation is also proposed to combine SDSC with MSE, improving stability and preserving amplitude where necessary. Experiments on forecasting and classification benchmarks demonstrate that SDSC-based pre-training achieves comparable or improved performance over MSE, particularly in in-domain and low-resource scenarios. The results suggest that structural fidelity in signal representations enhances the semantic representation quality, supporting the consideration of structure-aware metrics as viable alternatives to conventional distance-based methods.
@@ Introduction
To address these issues, a structure-aware metric, SDSC is introduced. The SDSC is derived from the DSC (Dice 1945; Sørensen 1948) widely used in semantic segmentation tasks and quantifies the structural agreement between two temporal signals. The metric is bounded in [0, 1], reduces sensitivity to amplitude variation, and reflects alignment in waveform structure. In addition, the metric is implemented as a loss by subtracting values from 1 and a differentiable approximation of the Heaviside function is applied to enable gradient-based optimization. Lastly, a hybrid loss that combines SDSC and MSE is introduced to enhance stability during training. Experiments on forecasting and classification benchmarks show that SDSC-based pre-training achieves comparable or improved performance relative to MSE, particularly when encoders are fixed. Performance gains are observed in in-domain and low-resource settings, where structural fidelity in the representations contributes to downstream accuracy.
@@
Although these works contributed significantly to the progress of TSM, most focused on architectural innovations or adopted contrastive strategies, which can be limiting to learning structural representations. To address this, we introduce a structure-aware metric tailored to time series data. SimMTM (Dong et al. 2023) is selected as our experimental baseline for its simplicity, strong empirical performance, and consistent application of traditional metrics, providing a reliable foundation to assess the impact of our proposed SDSC.
@@
Forecasting
  Pre-Model MSE SDSC Hybrid
Metrics MSE↓ MAE↓ MSE↓ MAE↓ MSE↓ MAE↓
ETTh1 0.380 0.408 0.379 0.406 0.382 0.406
ETTh2 0.304 0.350 0.306 0.352 0.304 0.350
ETTm1 0.327 0.363 0.324 0.364 0.325 0.364
ETTm2 0.185 0.275 0.191 0.278 0.188 0.276
Weather 0.169 0.213 0.168 0.213 0.169 0.215
Electricity 0.200 0.291 0.200 0.293 0.198 0.288
Traffic 0.497 0.312 0.492 0.309 0.494 0.315
Avg 0.295 0.316 0.294 0.316 0.294 0.316
Table 4: Forecasting Finetuning Example

### v2

@@ Abstract
We propose the Signal Dice Similarity Coefficient (SDSC), a novel structure-aware metric function for time series self-supervised representation learning. Most Self-Supervised Learning (SSL) methods for signals commonly adopt distance-based objectives such as mean squared error (MSE), which are sensitive to amplitude, invariant to waveform polarity, and unbounded in scale. These properties hinder semantic alignment and reduce interpretability. To the best of our knowledge, SDSC is the first similarity metric explicitly designed to capture structural alignment in temporal signals. SDSC addresses this by quantifying structural agreement between temporal signals based on the intersection of signed amplitudes, derived from the Dice Similarity Coefficient (DSC).Although SDSC is defined as a structure-aware metric, it can be used as a loss by subtracting from 1 and applying a differentiable approximation of the Heaviside function for gradient-based optimization. A hybrid loss formulation is also proposed to combine SDSC with MSE, improving stability and preserving amplitude where necessary. Experiments on forecasting and classification benchmarks demonstrate that SDSC-based pre-training achieves comparable or improved performance over MSE, particularly in in-domain and low-resource scenarios. The results suggest that structural fidelity in signal representations enhances the semantic representation quality, supporting the reconsideration of structure-aware metrics as viable alternatives to conventional distance-based methods.
@@ Introduction
To address these issues, a novel structure-aware metric, SDSC is introduced. The SDSC is inspired by the DSC (Dice 1945; Sørensen 1948) widely used in semantic segmentation tasks and quantifies the structural agreement between two temporal signals. The metric is bounded in [0, 1], reduces sensitivity to amplitude variation, and reflects alignment in waveform structure. In addition, the metric is implemented as a loss by subtracting values from 1 and a differentiable approximation of the Heaviside function is applied to enable gradient-based optimization. Lastly, a hybrid loss that combines SDSC and MSE is introduced to enhance stability during training. Experiments on forecasting and classification benchmarks show that SDSC-based pre-training achieves comparable or improved performance relative to MSE, particularly when encoders are fixed. Performance gains are observed in in-domain and low-resource settings, where structural fidelity in the representations contributes to downstream accuracy.
@@
Although these works contributed significantly to the progress of TSM, most focused on architectural innovations or adopted contrastive strategies, which can be limiting to learning structural representations. To the best of our knowledge, no prior work has proposed a structure-aware metric for signal representation learning. To address this, we introduce a structure-aware metric tailored to time-series data. SimMTM (Dong et al. 2023) is selected as our experimental baseline for its simplicity, strong empirical performance, and consistent application of traditional metrics, providing a reliable foundation to assess the impact of our proposed SDSC.
@@
Experiments
Gradient Sensitivity Analysis
In this section, we measure the gradient norm with respect to different types of signal perturbations to assess how each loss function responds to structural variations. The analysis is based on the same toy cases as in Table 1, including the jitter case.
  Example MSE MAE SDSC
Inverted 0.0894 0.0316 0.0
0.5x Scaled 0.0223 0.0316 0.0442
2x Scaled 0.0447 0.0316 0.0110
Zero 0.0447 0.0316 0.0
Noise Sample 0.0194 0.0316 0.0237
Shifted 0.0632 0.0316 0.0075
Jittered 0.0032 0.0316 0.0248
Table 2: Gradient Sensitivity Table
  Example $\alpha=1$ $\alpha=10$ $\alpha=100$
Inverted 0.0091 0.0082 0.0047
0.5x Scaled 0.0289 0.0437 0.0436
2x Scaled 0.0062 0.0102 0.0102
Zero 0.0 0.0 0.0
Noise Sample 0.0152 0.0228 0.0237
Shifted 0.0074 0.0087 0.0076
Jittered 0.0165 0.0228 0.0242
Table 3: Gradient Sensitivity with alpha Table
Tables 2 and 3 present the gradient norms for MSE, MAE, and SDSC, and the effect of varying the $\alpha$ parameter in SDSC, respectively.
MSE exhibits significant gradient changes under amplitude perturbations, but it fails to respond meaningfully to minor structural variations, such as jitter. In contrast, SDSC yields low gradients for structure-preserving signals (e.g., shifted), while responding with larger gradients when local waveform patterns are distorted (e.g., jittered). Although gradient vanishing may occur in some structure-breaking cases, SDSC remains robust to amplitude shifts.
As $\alpha$ increases, the behavior of the approximated SDSC becomes more similar to that of the original formulation. The results suggest that $\alpha=10$ is sufficient for a close approximation in practice.
@@
Forecasting
  Pre-Model MSE SDSC Hybrid
Metrics MSE↓ MAE↓ MSE↓ MAE↓ MSE↓ MAE↓
ETTh1 0.380 0.408 0.379 0.406 0.382 0.406
ETTh2 0.304 0.350 0.306 0.352 0.304 0.350
ETTm1 0.327 0.363 0.324 0.364 0.325 0.364
ETTm2 0.185 0.275 0.191 0.278 0.188 0.276
Weather 0.169 0.213 0.168 0.213 0.169 0.215
Electricity 0.200 0.291 0.200 0.293 0.198 0.288
Traffic 0.497 0.312 0.492 0.309 0.494 0.315
Avg 0.295 0.316 0.294 0.316 0.294 0.316
Table 7: Forecasting Finetuning Example